OpenStudy (anonymous):

Create a proportion using the form %/100 = is/of to represent the following word problems. Seven is what percent of 30? ?/? = ?/? What is 60% of 20? ?/? = ?/? 15 is 1 1/2 of what number? ?/? = ?/?

OpenStudy (texaschic101):

x/100 = 7/30 60/100 = x/20 not sure about this last one... 1.5/100 = 15/x

OpenStudy (anonymous):

you were right thank you so much! :)